# Distribution estimation via Flow Matching with Lipschitz guarantees

### v1

@@ 7.4 Additional proofs of Section˜7 and helper results
Thus
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$ $\lesssim\log(n)^{6}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}p_{t}(x)\;\mathrm{d}x$
$\quad+\log(n)^{4}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|x|^{2}p_{t}(x)\;\mathrm{d}x.$
@@
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{dy}$ $=\frac{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y\;\mathrm{d}x}$
$\leq e^{2L}\frac{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y}{\int\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y\;\mathrm{d}x}.$
@@
$\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y$ $=\exp(-\frac{|x|^{2}}{2\sigma_{t}^{2}})\int\exp(-\frac{1}{2}\big{(}1+\frac{t^{2}}{\sigma_{t}^{2}}\big{)}|y|^{2}+\langle\frac{t}{\sigma_{t}^{2}}x,y\rangle)\;\mathrm{d}y$
$=\exp(-\frac{|x|^{2}}{2\sigma_{t}^{2}}(2\pi)^{d/2}\big{(}\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\big{)}^{d/2}\exp(\frac{|x|^{2}}{2}\frac{t^{2}}{(t^{2}+\sigma_{t}^{2})}\sigma_{t}^{2}).$
@@
$p_{t}(x)$ $\leq e^{2L}\frac{\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})}{\int\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})\;\mathrm{d}x}=\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2}).$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}$ $p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|x|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$
$\leq\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}|x|^{2}\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})\;\mathrm{d}x\;\mathrm{d}t$
$=2\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\Big{(}\int_{\log(n)}^{\infty}x_{1}^{2}\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}x_{1}^{2})\;\mathrm{d}x_{1}\Big{)}^{d}\;\mathrm{d}t$
$\leq 2\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\Big{(}\log(n)(\sigma_{t}^{2}+t^{2})\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}\log(n)^{2})\Big{)}^{d}\;\mathrm{d}t,$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus[-\log(n),\log(n)]^{d}}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$ $\lesssim\log(n)^{d}\exp(-\frac{1}{4}\log^{2}(n)).$
Overall, we need to choose $n$ big enough such that
$\log(n)^{6+d}\exp(-\frac{1}{4}\log^{2}(n))\leq n^{-\frac{1}{2}}\quad\Longleftarrow\quad n\geq e^{4(\frac{1}{2}+6+d)},$
where again the implication stems from a loose upper bound using $\log(n)^{6+d}\leq n^{6+d}$.
Proof of Lemma˜7.6.
@@
$|v_{t}(x|y)|=\Big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x+(\sigma_{t}-t)y)\Big{|}\leq\Big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big{|}(|x|+|y|)\leq|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}|(|x|+\sqrt{d}\log(n)).$
@@
$\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}(\log(n)^{3}+\log(n)\int_{0}^{1}\Big{|}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big{|}\int|x|p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t)$
$\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}\log(n)^{4}.$
The last inequality follows from Jensens inequality and the same argument as in (34). This concludes the proof. ∎
Proof of Lemma˜7.7.
@@
where
$q_{t,x}\propto\exp(-\frac{\|x-ty\|^{2}}{2\sigma_{t}^{2}}-\frac{\|y\|^{2}}{2}-a(y))\propto\exp(\langle\eta,Y\rangle)h(y),$
where $\eta=\frac{t}{\sigma_{t}^{2}}x$ and $h(y)=\exp(-(\frac{t^{2}}{2\sigma_{t}^{2}}+\frac{1}{2})\|y\|^{2}-a(y))$. Define
$A(\eta)\coloneqq\log\int\exp(\langle\eta,Y\rangle)h(y)\;\mathrm{d}y.$
In context of exponential families, $A$ is typically called the log-partition function. Then
$\frac{\partial}{\partial\eta_{j}}A(\eta)=\mathbb{E}_{q_{t,x}}[Y_{j}].$
Thus,
@@
Further define the cumulant generating function as
$K(\lambda)=\log\big{(}\mathbb{E}[\exp(\langle\lambda,Y\rangle)]\big{)}.$
Then
@@
To control the moments up to order $s+1$, we need a stronger argument than the Poincaré inequality used for the variance bound.
Definition 7.9.
The standard Gaussian satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $\lambda=1$ (Boucheron et al., 2013, Theorem 5.4). Analogously to Poincaré inequality, its easy to see that $\mathcal{N}(\mu,\sigma^{2}I_{d})$ satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $\lambda=\sigma^{-2}$. Similar to the beginning of the proof of Theorem˜4.6, we conclude that the distribution with density (20) satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $(1+\frac{t^{2}}{\sigma_{t}^{2}})$. By (Ledoux, 2001, Lemma 1.2) we known that a bounded perturbation will again only impact the $\log$-Sobolev constant by an exponential term. Thus, we can bound the $\log$-Sobolev constant of the distribution of $Y^{x,t}$ by
$\lambda\geq e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}}).$ (43)
Now we can use Herbst’s argument (along the lines of Ledoux (1999, p.148), note that our function is so simple that we can drop the assumption of a bounded function) to show that $Y^{x,t}_{i}-\mathbb{E}[Y^{x,t}_{i}]$ is subgaussian. We omit the dependency on $x$ and $t$ in the following. Set $f(y)=e^{\frac{\vartheta y_{i}}{2}}$ for a $\vartheta>0$ in (42). Define $H(\vartheta)\coloneqq\mathbb{E}[e^{\vartheta Y_{i}}]$. Then using the dominated convergence theorem
$\operatorname{Ent}(f^{2})$ $=\mathbb{E}[\vartheta Y_{i}e^{\vartheta Y_{i}}]-\mathbb{E}[e^{\vartheta Y_{i}}]\log\big{(}\mathbb{E}[e^{\vartheta Y_{i}}]\big{)}=\vartheta H^{\prime}(\vartheta)-H(\vartheta)\log(H(\vartheta)).$
Further
$\frac{2}{\lambda}\mathbb{E}[|\nabla f(Y)|^{2}]$ $=\frac{\vartheta^{2}}{2\lambda}\mathbb{E}[e^{\vartheta Y_{i}}]=\frac{\vartheta^{2}\lambda}{2}H(\vartheta).$
Thus by (42)
$\vartheta H^{\prime}(\vartheta)-H(\vartheta)\log(H(\vartheta))\leq\frac{\vartheta^{2}}{2}H(\vartheta)\quad\Longleftrightarrow\quad\frac{1}{\vartheta}\frac{H^{\prime}(\vartheta)}{H(\vartheta)}-\frac{\log(H(\vartheta))}{\vartheta^{2}}\leq\frac{1}{2\lambda}.$
Now define $K(\vartheta)\coloneqq\frac{1}{\vartheta}\log(H(\vartheta))$. Then
$K^{\prime}(\vartheta)=\frac{H^{\prime}(\vartheta)}{H(\vartheta)}-\frac{\log(H(\vartheta))}{\vartheta^{2}}\leq\frac{1}{2\lambda}.$
As
$\lim_{\vartheta\rightarrow 0}K(\vartheta)=\lim_{\vartheta\rightarrow 0}\frac{\mathbb{E}[e^{\vartheta Y_{i}}]}{\vartheta}=\lim_{\vartheta\rightarrow 0}\frac{\mathbb{E}[Y_{i}e^{\vartheta Y_{i}}]}{\mathbb{E}[e^{\vartheta Y_{i}}]}=\lim_{\vartheta\rightarrow 0}\frac{M^{\prime}(\vartheta)}{M(\vartheta)}=\mathbb{E}[Y_{i}],$
where $M$ is the moment generating function, we can extend the definition of $K$ to
$K(\vartheta)\coloneqq\begin{cases}\frac{1}{\vartheta}\log(H(\vartheta)),&\vartheta>0,\\ \mathbb{E}[Y_{i}],&\vartheta=0.\end{cases}$
Hence
$K(\vartheta)=K(0)+\int_{0}^{\vartheta}K^{\prime}(u)\;\mathrm{d}u\leq\mathbb{E}[Y_{i}]+\frac{\vartheta}{2\lambda}.$
Going back to the function $H$, we get
$\frac{1}{\vartheta}\log(H(\vartheta))\leq\mathbb{E}[Y_{i}]+{\frac{1}{2\lambda}}\quad\Longleftrightarrow\quad H(\vartheta)\leq\exp(\vartheta\mathbb{E}[Y_{i}]+\frac{\vartheta^{2}}{2\lambda})\quad\Longleftrightarrow\quad\mathbb{E}[e^{\vartheta Y_{i}}]\leq\exp(\vartheta\mathbb{E}[Y_{i}]+\frac{\vartheta^{2}}{2\lambda}).$
Now Markov’s inequality yields
$\mathbb{P}(Y_{i}-\mathbb{E}[Y_{i}]\geq r)$ $=\mathbb{P}\big{(}\exp(\vartheta(Y_{i}-\mathbb{E}[Y_{i}]))\geq\exp(\vartheta r)\big{)}\leq\frac{\mathbb{E}[\exp(\vartheta(Y_{i}-\mathbb{E}[Y_{i}]))]}{\exp(\vartheta r)}$
$=\exp(-\vartheta r)\frac{\mathbb{E}[e^{\vartheta Y_{i}}]}{e^{\vartheta\mathbb{E}[Y_{i}]}}\leq\exp(-\vartheta r)\frac{\exp(\vartheta\mathbb{E}[Y_{i}]+\frac{\vartheta^{2}}{2\lambda})}{e^{\vartheta\mathbb{E}[Y_{i}]}}=e^{\frac{\vartheta^{2}}{2\lambda}-\vartheta r}.$
Setting $f(x)=e^{-\frac{\vartheta x_{i}}{2}}$, we obtain analogously
$\mathbb{P}(Y_{i}-\mathbb{E}[Y_{i}]\leq-r)\leq e^{\frac{\vartheta^{2}}{2\lambda}-\vartheta r}.$
Thus a union bound argument and setting $\vartheta=r\lambda$ leads to
$\mathbb{P}(|Y_{i}-\mathbb{E}[Y_{i}]|\geq r)\leq 2e^{-\frac{r^{2}\lambda}{2}}.$
Now we can bound the $s+1$-th moment along the lines of (Vershynin, 2018, Proposition 2.5.2) via
@@
$\leq 2(s+1)\int_{0}^{\infty}e^{-\frac{r^{2}\lambda}{2}}r^{s}\;\mathrm{d}r=2^{\frac{s+1}{2}}\lambda^{-\frac{s+1}{2}}\Gamma(\frac{s+1}{2}).$
Taking the $s+1$-th root leads to the bound
$\mathbb{E}[|Y_{i}-\mathbb{E}[Y_{i}]|^{s+1}]^{\frac{1}{s+1}}$ $\lesssim\lambda^{-\frac{1}{2}}.$
Plugging in the $\log$ Sobolev constant of $Y$ from (43), we can control all moments via
@@
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)\lesssim\mathds{1}_{s=1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+(1-\frac{\sigma_{t}^{\prime}t}{\sigma_{t}})\big{(}\frac{t}{\sigma_{t}^{2}}\big{)}^{s}\min\Big{(}1,\frac{\sigma_{t}^{s+1}}{t^{s+1}}\Big{)}.$
Minimizing in $\sigma_{t}$ and using $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min})$ leads to
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)\lesssim\log(\sigma_{\min})\sigma_{\min}^{-s+1}.$
Proof of Lemma˜7.8.
@@
$q_{x,t}(y)\coloneqq\frac{\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))}{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y},\quad w_{x,t}(y)\coloneqq\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y)).$
@@
$\frac{\partial}{\partial t}v^{j}_{t}(x)$ $=\frac{\partial}{\partial t}\mathbb{E}_{Y\sim q_{x,t}}[v_{t}(x|Y)_{j}]$ (44)
$=\mathbb{E}_{Y\sim q_{x,t}}\Big{[}\frac{\partial}{\partial t}v_{t}(x|Y)_{j}\Big{]}+\mathbb{E}_{Y\sim q_{x,t}}\Big{[}v_{t}(x|Y)_{j}\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big{]}$
$\quad-\mathbb{E}_{Y\sim q_{x,t}}\Big{[}v_{t}(x|Y)_{j}\Big{]}\mathbb{E}_{Y\sim q_{x,t}}\Big{[}\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big{]}$
$=\mathbb{E}_{Y\sim q_{x,t}}\Big{[}\frac{\partial}{\partial t}v_{t}(x|Y)_{j}\Big{]}+\operatorname{Cov}\Big{(}v_{t}(x|Y)_{j},\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big{)}.$
@@
$v_{t}(x|y)_{j}$ $=\log(\sigma_{\min})(x_{j}-ty_{j})+y_{j},$ (45)
$\frac{\partial}{\partial t}v_{t}(x|y)_{j}$ $=\log(\sigma_{\min})y_{j},$
@@
$\frac{\partial}{\partial t}\log(w_{x,t}(Y))$ $=-\frac{\partial}{\partial t}\frac{1}{2}(|x-ty|^{2})\sigma_{t}^{-2}$
$=(\langle x,y\rangle-t|y|^{2})\sigma_{t}^{-2}+\frac{1}{2}|x-ty|^{2}\cdot 2\log(\sigma_{\min})\sigma_{t}^{-2}.$ (46)
@@
$\operatorname{Cov}((\frac{\partial}{\partial t}$ $\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y)))$
$=\mathbb{E}\Big{[}\Big{(}\frac{\partial}{\partial t}\log(w_{x,t}(Y))-\mathbb{E}\Big{[}\frac{\partial}{\partial t}\log(w_{x,t}(Y))\big{]}\Big{)}^{k}\Big{]}$
@@
$\qquad\qquad+(t^{2}2\log(\sigma_{\min})-t)|Y|^{2}-\mathbb{E}[(t^{2}2\log(\sigma_{\min})-t)|Y^{2}|]\Big{)}^{k}\Big{]}$
@@
where all expectations are taken with respect to $Y\sim q_{x,t}$. From the proof of Lemma˜7.7 we know that
$\mathbb{E}\Big{[}\big{(}Y_{i}-\mathbb{E}[Y_{i}]\big{)}^{k}\Big{]}\lesssim\big{(}\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\big{)}^{\frac{k}{2}}\leq\min(\frac{\sigma_{t}^{2}}{t^{2}},1)^{\frac{k}{2}}=\min(\frac{\sigma_{t}^{k}}{t^{k}},1).$
By Vershynin (2018, Lemma 2.7.6) combined with Vershynin (2018, Lemma 2.7.1) we know that if $Y_{i}-\mathbb{E}[Y_{i}]$ is subgaussian, which was shown in Lemma˜7.7, then $(Y_{i}-\mathbb{E}[Y_{i}])^{2}$ is subexponential with the same order of subgaussian and subexponential norm. For a definition of a subexponential random variables, the subexponential norm and the subgaussian norm, we refer to Vershynin (2018, Section 2.7). Note that all the constants in Vershynin (2018, Proposition 2.7.1) only differ by a constant.

### v2

@@ Appendix A Additional proofs of Section˜6 and helper results
Thus, setting $I=[-\log(n),\log(n)]^{d}$
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}|v_{t}(x)-\tilde{v}_{t}(x)|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t$ $\lesssim\log(n)^{6}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}p_{t}(x)\;\mathrm{d}x+\log(n)^{4}\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}|x|^{2}p_{t}(x)\;\mathrm{d}x.$
@@
$p_{t}(x)=\int p_{t}(x|y)p^{*}(y)\;\mathrm{dy}=\frac{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y}{\int\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y\;\mathrm{d}x}e^{2L}\frac{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y}{\int\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2})\;\mathrm{d}y\;\mathrm{d}x}.$
@@
$\int\exp\Big(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}\Big)\;\mathrm{d}y=(2\pi)^{d/2}\Big(\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\Big)^{d/2}\exp\Big(-\frac{|x|^{2}}{2\left(t^{2}+\sigma_{t}^{2}\right)}\Big).$
@@
$p_{t}(x)\leq e^{2L}\frac{\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})}{\int\exp(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2})\;\mathrm{d}x}=\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\exp\Big(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2}\Big).$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}$ $p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t\leq\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}|x|^{2}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t\leq\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\int_{\mathbb{R}^{d}\setminus I}|x|^{2}\exp\Big(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}|x|^{2}\Big)\;\mathrm{d}x\;\mathrm{d}t$
$=2\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\Big(\int_{\log(n)}^{\infty}x_{1}^{2}\exp\Big(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}x_{1}^{2}\Big)\;\mathrm{d}x_{1}\Big)^{d}\;\mathrm{d}t$
$\leq 2\int_{0}^{1}\frac{e^{2L}}{(2\pi(\sigma_{t}^{2}+t^{2}))^{\frac{d}{2}}}\Big(\log(n)(\sigma_{t}^{2}+t^{2})\exp\Big(-\frac{1}{2}\frac{1}{\sigma_{t}^{2}+t^{2}}\log(n)^{2}\Big)\Big)^{d}\;\mathrm{d}t,$
@@
$\int_{0}^{1}\int_{\mathbb{R}^{d}\setminus I}p_{t}(x)\;\mathrm{d}x\;\mathrm{d}t\lesssim\log(n)^{d}\exp\Big(-\frac{1}{4}\log^{2}(n)\Big).$
Overall, we need to choose $n$ big enough such that $\log(n)^{6+d}\exp(-\frac{1}{4}\log^{2}(n))\leq n^{-\frac{1}{2}}$ which is implied by $n\geq e^{4(\frac{1}{2}+6+d)},$ where again the implication stems from a loose upper bound using $\log(n)^{6+d}\leq n^{6+d}$. ∎
Proof of Lemma˜6.7.
@@
$|v_{t}(x|y)|=\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}(x+(\sigma_{t}-t)y)\Big|\leq\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|(|x|+|y|)\leq\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|(|x|+\sqrt{d}\log(n)).$
@@
$\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}(\log(n)^{3}+\log(n)\int_{0}^{1}\Big|\frac{\sigma_{t}^{\prime}}{\sigma_{t}}\Big|\int|x|p_{t}(x|y)\;\mathrm{d}x\;\mathrm{d}t)\lesssim\|\bar{v}^{1}-\bar{v}^{2}\|_{\infty}\log(n)^{4}.$
The last inequality follows from Jensens inequality and the properties of the Gaussian. ∎
Proof of Lemma˜6.8.
@@
where $q_{t,x}\propto\propto\exp(\langle\eta,Y\rangle)h(y),$ with $\eta=\frac{t}{\sigma_{t}^{2}}x$ and $h(y)=\exp(-(\frac{t^{2}}{2\sigma_{t}^{2}}+\frac{1}{2})\|y\|^{2}-a(y))$. Define $A(\eta)\coloneqq\log\int\exp(\langle\eta,Y\rangle)h(y)\;\mathrm{d}y.$ In context of exponential families, $A$ is typically called the log-partition function. Then $\frac{\partial}{\partial\eta_{j}}A(\eta)=\mathbb{E}_{q_{t,x}}[Y_{j}].$ Thus,
@@
Further define the cumulant generating function as $K(\lambda)\coloneqq\log\big(\mathbb{E}[\exp(\langle\lambda,Y\rangle)]\big).$ Then
@@
To control the moments up to order $s+1$, we need a stronger argument than the Poincaré inequality used for the variance bound. Thus, we recall Definition˜6.4, the definition of the $\log$-Sobolev inequality. Similar to the beginning of the proof of Proposition˜3.6, we conclude that the distribution with density (21) satisfies the $\log$-Sobolev inequality with $\log$-Sobolev constant $(1+\frac{t^{2}}{\sigma_{t}^{2}})$. By Ledoux (2001, Lemma 1.2) we known that a bounded perturbation will again only impact the $\log$-Sobolev constant by an exponential term. Thus, we can bound the $\log$-Sobolev constant of the distribution of $Y^{x,t}$ by $\lambda\geq e^{-4L}(1+\frac{t^{2}}{\sigma_{t}^{2}}).$
Now we can use Herbst’s argument (along the lines of Ledoux (1999, p.148), note that our function is so simple that we can drop the assumption of a bounded function) to conclude that $Y^{x,t}_{i}-\mathbb{E}[Y^{x,t}_{i}]$ is subgaussian, i.e.
$\mathbb{P}(|Y_{i}-\mathbb{E}[Y_{i}]|\geq r)\leq 2e^{-\frac{r^{2}\lambda}{2}}.$ Then we can bound the $s+1$-th moment along the lines of Vershynin (2018, Proposition 2.5.2) via
@@
$\leq 2(s+1)\int_{0}^{\infty}e^{-\frac{r^{2}\lambda}{2}}r^{s}\;\mathrm{d}r=2^{\frac{s+1}{2}}\lambda^{-\frac{s+1}{2}}\Gamma\Big(\frac{s+1}{2}\Big).$
Taking the $s+1$-th root leads to the bound $\mathbb{E}[|Y_{i}-\mathbb{E}[Y_{i}]|^{s+1}]^{\frac{1}{s+1}}\lesssim\lambda^{-\frac{1}{2}}.$ Plugging in the $\log$ Sobolev constant of $Y$, we can control all moments via
@@
$\frac{\partial^{s}}{\partial x_{i_{1}},...,\partial x_{i_{s}}}v^{j}_{t}(x)\lesssim\mathds{1}_{s=1}\frac{\sigma_{t}^{\prime}}{\sigma_{t}}+\Big(1-\frac{\sigma_{t}^{\prime}t}{\sigma_{t}}\Big)\Big(\frac{t}{\sigma_{t}^{2}}\Big)^{s}\min\Big(1,\frac{\sigma_{t}^{s+1}}{t^{s+1}}\Big).$
Minimizing in $\sigma_{t}$ and using $\frac{\sigma_{t}^{\prime}}{\sigma_{t}}=\log(\sigma_{\min})$ leads to the final bound. ∎
Proof of Lemma˜6.9.
@@
$q_{x,t}(y)\coloneqq\frac{\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))}{\int\exp(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y))\;\mathrm{d}y},\quad w_{x,t}(y)\coloneqq\exp\Big(-\frac{|x-ty|^{2}}{2\sigma_{t}^{2}}-\frac{|y|^{2}}{2}-a(y)\Big).$
@@
$\frac{\partial}{\partial t}v^{j}_{t}(x)=\frac{\partial}{\partial t}\mathbb{E}_{Y\sim q_{x,t}}[v_{t}(x|Y)_{j}]=\mathbb{E}_{Y\sim q_{x,t}}\Big[\frac{\partial}{\partial t}v_{t}(x|Y)_{j}\Big]+\operatorname{Cov}\Big(v_{t}(x|Y)_{j},\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big).$ (42)
@@
$v_{t}(x|y)_{j}=\log(\sigma_{\min})(x_{j}-ty_{j})+y_{j},\quad\frac{\partial}{\partial t}v_{t}(x|y)_{j}=\log(\sigma_{\min})y_{j},$ (43)
@@
$\frac{\partial}{\partial t}\log(w_{x,t}(Y))=-\frac{\partial}{\partial t}\frac{1}{2}(|x-ty|^{2})\sigma_{t}^{-2}=(\langle x,y\rangle-t|y|^{2})\sigma_{t}^{-2}+\frac{1}{2}|x-ty|^{2}\cdot 2\log(\sigma_{\min})\sigma_{t}^{-2}.$ (44)
@@
$\operatorname{Cov}((\frac{\partial}{\partial t}\log(w_{x,t}(Y)))^{k-1},\frac{\partial}{\partial t}\log(w_{x,t}(Y)))=\mathbb{E}\Big[\Big(\frac{\partial}{\partial t}\log(w_{x,t}(Y))-\mathbb{E}\Big[\frac{\partial}{\partial t}\log(w_{x,t}(Y))\Big]\Big)^{k}\Big]$
@@
$\quad+(t^{2}2\log(\sigma_{\min})-t)|Y|^{2}-\mathbb{E}[(t^{2}2\log(\sigma_{\min})-t)|Y^{2}|]\Big)^{k}\Big]$
@@
where all expectations are taken with respect to $Y\sim q_{x,t}$. From the proof of Lemma˜6.8 we know that
$\mathbb{E}\Big[\big(Y_{i}-\mathbb{E}[Y_{i}]\big)^{k}\Big]\lesssim\Big(\frac{\sigma_{t}^{2}}{t^{2}+\sigma_{t}^{2}}\Big)^{\frac{k}{2}}\leq\min\Big(\frac{\sigma_{t}^{2}}{t^{2}},1\Big)^{\frac{k}{2}}=\min\Big(\frac{\sigma_{t}^{k}}{t^{k}},1\Big).$
By Vershynin (2018, Lemma 2.7.6) combined with Vershynin (2018, Lemma 2.7.1) we know that if $Y_{i}-\mathbb{E}[Y_{i}]$ is subgaussian, which was shown in Lemma˜6.8, then $(Y_{i}-\mathbb{E}[Y_{i}])^{2}$ is subexponential with the same order of subgaussian and subexponential norm. For a definition of a subexponential random variables, the subexponential norm and the subgaussian norm, we refer to Vershynin (2018, Section 2.7). Note that all the constants in Vershynin (2018, Proposition 2.7.1) only differ by a constant.